OpenStudy (ams98):

Explain these to me so I don't get the same type question wrong again!!! Complex Numbers! MEDAL!!

OpenStudy (ams98):

@GABI_J

OpenStudy (ams98):

(−3 + 2i) • (2 + i)

OpenStudy (ams98):

What do you mean??

OpenStudy (cggurumanjunath):

FOIL METHOD.

OpenStudy (cggurumanjunath):

http://www.mathwords.com/f/foil_method.htm

OpenStudy (ciel21):

if i explain 9/10 chances i wont get a medal :'c

OpenStudy (anonymous):

Yes, Foil. \[(a+b)(c+d)\] \[ac+ad+bd+bc\]

OpenStudy (ams98):

Okay let me try one sec!

OpenStudy (anonymous):

\[(−3 + 2i) • (2 + i)\] Recalling)

OpenStudy (ams98):

Oh gawsh i still probably got it wrong as I suck at math but I got -8-i

OpenStudy (anonymous):

Almost, look over your work again!

OpenStudy (anonymous):

One part is wrong the other is right? Which one is wrong -8 or -i what do you think?

OpenStudy (ams98):

-8+i??

OpenStudy (ams98):

I did wrong signs

OpenStudy (anonymous):

Yes,

OpenStudy (anonymous):

you are right now

OpenStudy (ams98):

Okay one more! And thanks so much!!

OpenStudy (anonymous):

Ask away!

OpenStudy (ams98):

(1 + 2i) • (5 − 3i)

OpenStudy (ams98):

let me try... :)

OpenStudy (anonymous):

you can do it!

OpenStudy (ams98):

11+7i?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Can I give you a harder example?

OpenStudy (ams98):

Ahh thanks so much! Your a life saver!!

OpenStudy (ams98):

and sure!

OpenStudy (anonymous):

OK, ready? \[\frac{5+i}{2-i}\] This is also trig, but a little harder Hint to think_ you can have a radical in a denominator Any ideas?

OpenStudy (ams98):

hmmm give me a hint!

OpenStudy (anonymous):

Multiply the equation by \[\frac{2+i}{2+i}\]. this is what I refer to as a magical 1.

OpenStudy (anonymous):

I'll right it out for you, and you multiply, Ok?

OpenStudy (ams98):

always? OK

OpenStudy (anonymous):

\[\frac{(5+i)(2+i)}{(2-i)(2+i)}\] Right? Do this one for me real quick.

OpenStudy (ams98):

What do you do with the I's? Sqaure them? But if you square them arent they just like a -1?

OpenStudy (anonymous):

i^2 is exactly -1

OpenStudy (anonymous):

So write it out for me please....

OpenStudy (ams98):

Okay! So I am ushally wrong but....\[\frac{ 10-1 }{ 4-1 }=\frac{ 9 }{ 3 }= 3\]

OpenStudy (anonymous):

No, ain't right!

OpenStudy (anonymous):

You forgot about the middle term on top!

OpenStudy (anonymous):

\[(5+i)(2+i)\] what is the middle term?

OpenStudy (ams98):

Im lost!! I try re-doing it and I got something like this :(\[\frac{ 9+10i }{ 3+0 }\]

OpenStudy (anonymous):

Nope!

OpenStudy (anonymous):

9 is right, what is wrong?

OpenStudy (anonymous):

You got confused it is not 2*5i it is 2i+5i Don't you see that?

OpenStudy (anonymous):

\[(a+b)(c+d)_\] \[a(c+d)+b(c+d)_\] \[ac+ad+bc+bd_\]

OpenStudy (ams98):

wait so just 10 lmao?

OpenStudy (anonymous):

No, it is 9+7i b/c the number of i s is 2+5 see (I mean if you foil, step by step)

OpenStudy (ams98):

geeez I'm terrible at math xD

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

\[Understand?\]

OpenStudy (ams98):

kind of sort of!